Combine. Be sure to simplify your answer whenever possible
5/3 + 5/6

Answer by stanbon(75887)

5/3 + 5/6
lcd = 6
= 10/6 + 5/6
= (10+5)/6
= 15/6
= 5/2
============
Cheers,
Stan H.